You just opened a brokerage account, depositing $2,000. You expect the account to earn an interest rate of 8.84%. You also plan on depositing $2,000 at the end of years 5 through 10. What will be the value of the account at the end of 20 years, assuming you earn your expected rate of return? Please show the steps to use on a financial calculator.
Answer #1

Initial Deposit =2000
Rate =8.84%
Using financial calculator
N=20; I/Y =8.84%; PV=-2000; CPT FV =10884.30


PMT=2000
Number of years of deposit =6 years
PMT =2000
N=6;I/Y =8.84%; PMT =-2000; CPT FV =14986.05

To calculate FV of this amount
N=10;I/Y =8.84%; PV =-14986.05; CPT FV =34960.09

Value at end of 20 years =10.884.30+34960.09=45844.39